OpenStudy (anonymous):

How much would $220 invested at 10% interest compounded continuously be worth after 12 years? Round your answer to the nearest cent. Do not include units in your answer.

OpenStudy (anonymous):

\[220\times e^{1.2}\] and a calculator

OpenStudy (anonymous):

let me know where i can invest at 10% compounded continuously

OpenStudy (anonymous):

220 x 1,1^12 = 690.45

OpenStudy (phi):

Use this formula http://cs.selu.edu/~rbyrd/math/continuous/ sat filled in the details for you

OpenStudy (anonymous):

\[A(t)=P*e ^{rt}\]

OpenStudy (phi):

Yes, that is the formula. now we need to replace the letters with numbers what is P ? (they tell you) r is the rate. but r must be a decimal not a percent. change 10% to 0.1 to get r=0.1 what is t (time in years) ?

OpenStudy (anonymous):

Use williams answer if you want, but it's incorrect. He compounded just once annually, not continuously. Learn how to do this for yourself. You have all of the info you need just plug it into your calculator. P = principal = 220 e = 2.71 r = 10% t = 12 years A(t) is the answer you want. If you really want to know how this works and are curious why e shows up(it pops up in a lot of interesting places in math and physics) then look at this link. Very interesting stuff if you're willing to spend some time on it and actually learn it instead of just asking for the answers.